OpenStudy (anonymous):

how to convert 60km/h into m/s

OpenStudy (jamesj):

When you're new to this, I think it's easiest to do step by step 60 km/hr = 60,000 m/hr, because 1 km = 1,000 m = 60,000/60 m/min, because 1 hour = 60 minutes = 1,000 m/min = 1,000/60 m/sec, because 1 minute = 60 seconds = .... m/sec

OpenStudy (anonymous):

60 * 1000/3600

OpenStudy (ash2326):

You can multiply 60 Km/h by 5/18 to convert into m/s

OpenStudy (jamesj):

Yes ... now explain why your methods work. We don't like black boxes.

OpenStudy (anonymous):

oh 1KM = 1000m, and 1hr = 3600 seconds I hope that makes it clear

OpenStudy (anonymous):

I got 16.67 m/s.

OpenStudy (ash2326):

it's correct

OpenStudy (anonymous):

Thanks for your help!

OpenStudy (ash2326):

Kylie did you get why i multiplied by 5/18?

OpenStudy (anonymous):

No I didn't do it that way. Where did that come from?

OpenStudy (ash2326):

we have KM/hour 1 KM=1000 m 1 hour= 3600 seconds so 1 KM/hour= 1000m/3600 seconds or 1KM/hour= 10/36 m/sec or 1 km/hour= 5/18 m/sec so to convert km/hour into m/s , multiply by 5/18

OpenStudy (anonymous):

I got it. thanks I didn't break it all the way down.

OpenStudy (anonymous):

i think this way is conveniece \[60(km/h)\times(1h/3600s)\times(1000m/1km)\]